OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2ð). cos 4x - cos 2x = 0

OpenStudy (helder_edwin):

cos is a periodic function. the period is \(2\pi\), i.e. \[ \large \cos x=\cos(x+2k\pi) \] where k is an integer

OpenStudy (noelgreco):

What is the upper end of that interval?

OpenStudy (anonymous):

o 2pi/3 4pi/3

OpenStudy (anonymous):

Apply Cos C - Cos D formula

OpenStudy (anonymous):

cos C - Cos D = 2 sin(C+D)/2 Sin(D-C)/2

OpenStudy (anonymous):

so 0 pi/3 2pi/3 4pi/3 5pi/3

OpenStudy (anonymous):

so 2 Sin3x Sin(-x) =0 either Sin3x=0 or Sin x=0

OpenStudy (anonymous):

and use if Sin x=0 then x = n pi, where n =......-2,-1,0,1,2,3....

OpenStudy (anonymous):

now it's ok?

OpenStudy (anonymous):

im a lil confused what my anserw would be

OpenStudy (anonymous):

0 pi/3 2pi/3 4pi/3 5pi/3, pi, 2pi

OpenStudy (anonymous):

i think 2 pi is not in the interval...so remove 2 pi

OpenStudy (anonymous):

0 pi/3 2pi/3 4pi/3 5pi/3, pi

OpenStudy (anonymous):

now it's fine?

OpenStudy (anonymous):

either Sin3x=0 or Sin x=0 3x = n pi or x= n pi , where n =...-2,-1,0,1,2,3....

OpenStudy (anonymous):

oh ok